# WRF-GS: Wireless Radiation Field Reconstruction with 3D Gaussian Splatting

### v2

@@ VI-A Experiment Setup
VI-B Channel Estimation Accuracy
Fig. 13 depicts the prediction accuracy of the WRF-GS framework for the downlink CSI prediction task. The blue solid line represents the input ground truth uplink (GT Uplink) CSI, the orange solid line represents the ground truth downlink (GT Downlink) CSI, and the black dashed line is the predicted (Pred Downlink) CSI output by WRF-GS. It is seen that the predicted result is very close to the ground truth.
Figure 13: The comparison of the ground truth CSI and the predicted CSI using WRF-GS in terms of the channel amplitude and phase via different sub-carries.
To quantify the prediction accuracy, we use the SNR metric:
$\text{SNR}=-10\log_{10}\left({\left\|\mathrm{}C_{\text{pred}}-C_{\text{GT}}\right\|^{2}}/{\left\|C_{\text{GT}}\right\|^{2}}\right),$ (17)
where $C_{\text{pred}}$ is predicted downlink channel and $C_{\text{GT}}$ is ground truth downlink channel. The higher the SNR, the better the prediction accuracy. Fig. 14 shows the SNRs of five CSI prediction algorithms. Notably, WRF-GS achieves a median SNR of $22.98$ dB ($10$th percentile: $20.82$ dB, $90$th percentile: $24.33$ dB). The median SNRs of NeRF2, FIRE, R2F2, and OptML are $20.55$ dB, $15.29$ dB, $8.57$ dB, and $8.47$ dB, respectively. This demonstrates that WRF-GS improves the CSI prediction accuracy by efficiently modeling wireless channels.
Figure 14: The prediction accuracy of different methods in the downlink CSI prediction task. The figure shows the maximum, minimum, median, upper and lower quartile ranges of the SNR distribution for different methods, as well as the outliers. The higher the SNR, the better the prediction accuracy.
VII Conclusion
This paper introduced WRF-GS, a novel framework that combines 3D-GS and neural networks to enable fast and accurate wireless channel modeling. To adapt the 3D-GS technique from the optical domain to the RF domain, we designed three key modules for WRF-GS: a scenario representation network, a projection model, and an electromagnetic splatting technique. These modules allow WRF-GS to accurately capture the underlying physics of EM wave propagation. Remarkably, WRF-GS can synthesize new channel characteristics within milliseconds using a small number of training samples. Experimental results showed that WRF-GS outperforms existing methods in spatial spectrum reconstruction and excels in the CSI prediction task, achieving an improvement of over 2.43 dB. We envision that WRF-GS will enable various applications in future networks by leveraging its high-fidelity visualization, low sample complexity, and rapid rendering speed. The authors have provided public access to their code at https://github.com/wenchaozheng/WRF-GS.

### v4

@@ VII-B1 Experiment Setup
VII-B2 Channel Estimation Accuracy
Fig. 16 depicts the prediction accuracy of the WRF-GS and WRF-GS+ methods in the downlink CSI prediction task. Since the predicted CSI is a complex value, we measure the prediction accuracy on the amplitude and phase parts. We see that the predicted results are close to the ground truth. This means that the proposed methods can successfully predict the downlink CSI by extracting information from the uplink CSI based on the neural network architecture and 3D-GS.
Figure 17: The prediction accuracy of different methods in the downlink CSI prediction task. The figure shows the maximum, minimum, median, upper and lower quartile ranges of the CEA distribution for different methods, as well as the outliers. The higher the CEA, the better the prediction accuracy.
Next, we adopt the following metric to quantify the prediction accuracy and refer to it as CEA, representing the channel estimation accuracy in dB:
$\text{CEA}=-10\log_{10}\left(\frac{\left\|\mathrm{}C_{\text{pred}}-C_{\text{GT}}\right\|^{2}}{\left\|C_{\text{GT}}\right\|^{2}}\right),$ (20)
where $C_{\text{pred}}$ is predicted downlink channel and $C_{\text{GT}}$ is ground truth downlink channel. The higher the CEA, the better the prediction accuracy. Fig. 17 compares the prediction accuracy of the proposed methods and baselines. The results show that WRF-GS+ achieves a median CEA of $23.91$ dB ($10$th percentile: $20.85$ dB, $90$th percentile: $25.52$ dB), while WRF-GS achieves a median CAE of $22.98$ dB ($10$th percentile: $20.82$ dB, $90$th percentile: $24.33$ dB). In contrast, the median CEAs of NeRF2, FIRE, R2F2, and OptML are $20.55$ dB, $15.29$ dB, $8.57$ dB, and $8.47$ dB, respectively. The proposed methods consistently outperform the baselines. Notably, WRF-GS+ achieves the highest prediction accuracy, surpassing the best baseline NeRF2 by a significant margin of $3.36$ dB. This demonstrates that the superior capability of the proposed methods in efficiently and accurately modeling wireless channels.
VIII Conclusion
This paper first introduced WRF-GS, a novel framework for wireless channel modeling based on WRF reconstruction using neural network and 3D-GS. WRF-GS effectively captures the interactions between the environment and radio signals by refining the optical 3D-GS technique from three key aspects: scenario representation network, projection model, and electromagnetic splatting. To improve efficiency and accuracy, we further proposed WRF-GS+ by incorporating deformable 3D Gaussians and $\alpha$-blending to the scenario representation network and electromagnetic splatting modules, respectively. Experimental results showed that the proposed methods outperform the baselines for spatial spectrum synthesis. They can synthesize new channel characteristics within milliseconds using a small number of training samples. Moreover, WRF-GS+ achieves superior performance in both RSSI and CSI prediction tasks, surpassing existing methods by more than $0.7$ dB and $3.36$ dB, respectively. We envision that the proposed methods will enable various applications in future networks by leveraging their high-fidelity visualization, low sample complexity, and rapid rendering speed.